OpenStudy (unklerhaukus):

find the function whose transforms is given below. \[\frac{e^{-ap}}{p^2},\qquad a>0\]

OpenStudy (unklerhaukus):

\[ \begin{align*} f(t)&=\mathcal L ^{-1}\left\{\frac{e^{-ap}}{p^2}\right\},\quad a>0\\ &=\mathcal L ^{-1}\left\{e^{-ap}\times\frac1{p^2}\right\}\\ &= \end{align*}\]

OpenStudy (accessdenied):

Are you familiar with this theorem about convolution? \( \displaystyle \mathcal{L}^{-1} \left\{ F(s) G(s) \right\}(t) = \mathcal{L}^{-1} \{F(s)\}(t) * \mathcal{L}^{-1} \{G(s)\}(t) \) \( \displaystyle = (f * g)(t) \) \( \displaystyle = \int_{0}^{t} f(t - \tau) g(\tau) \; \textrm{d} \tau \) I found this in my PDF. If we can find the inverse Laplace transform of our two factors individually, then their convolution should give the inverse Laplace transform of their product.

OpenStudy (unklerhaukus):

i was trying to apply that but i got stuck

OpenStudy (accessdenied):

Hmm.. where did you get stuck?

OpenStudy (unklerhaukus):

Second Shift Theorem

OpenStudy (accessdenied):

Hmm... \( \mathcal{L} \{H(t-a) f(t-a) \}(s) = e^{-as} F(s) \) \( H(t-a) f(t-a) = \mathcal{L}^{-1} \{e^{-as} F(s) \} \) \(\displaystyle F(s) = \frac{1}{s^2} \) in this case. Looking at this, I think we simply need the inverse laplace transform of 1/s^2 to find f(t), and then sub t=t-a. I think that's possible w/o convolution... :)

OpenStudy (accessdenied):

From a table, I find the inverse laplace transform of 1/s^2 is f(t) = t. f(t-a) = (t-a) So the final answer would be... \( \displaystyle \mathcal{L}^{-1} \{\frac{e^{-as}}{s^2}\}(t) = (t-a) H(t-a) \)

OpenStudy (unklerhaukus):

\[\begin{align*} f(t)&=\mathcal L ^{-1}\left\{\frac{e^{-ap}}{p^2}\right\},\quad a>0\\ &=\mathcal L ^{-1}\left\{\frac{e^{-ap}}p\times\frac1{p}\right\}\\ &=\mathcal L ^{-1}\left\{\frac{e^{-ap}}p\right\}\times\mathcal L^{-1} \left\{\frac1{p}\right\}_{t\rightarrow t-a}\\ &=h(t-a)\times t\Big|_{t\rightarrow t-a}\\ &=h(t-a)\ (t-a)\\ \end{align*}\]??

OpenStudy (unklerhaukus):

im getting muddled up

OpenStudy (unklerhaukus):

\[\begin{align*} f(t)&=\mathcal L ^{-1}\left\{\frac{e^{-ap}}{p^2}\right\},\quad a>0\\ &=\mathcal L ^{-1}\left\{{e^{-ap}}\times\frac1{p^2}\right\}\\ &=h(t-a)\times\mathcal L^{-1} \left\{\frac1{p^2}\right\}_{t\rightarrow t-a}\\ &=h(t-a)\times t\Big|_{t\rightarrow t-a}\\ &=h(t-a)(t-a) \end{align*}\] ?

OpenStudy (accessdenied):

Yes, that is what I got as well. :)

OpenStudy (unklerhaukus):

ok i still dont understand this very well

OpenStudy (accessdenied):

I apologize for not being very much help with the intuition part here. I am not very familiar with it myself. ;(

OpenStudy (accessdenied):

My current understanding is that this theorem establishes that a factor of e^(-ap) on the laplace transform of a function translates to a shift of 'a' units to the right in the original function; or in reverse, that a shift of 'a'units translates to a factor of e^(-ap) in the transform.